Fermi National Laboratory
 

Accelerator Update for June 15, 2007 - June 18, 2007

 

Friday June 15

The day shift began with the Tevatron (TeV) doing studies, the Recycler maintaining a stash (124.2E10), with the Antiproton Source (Pbar) stacking (51.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam.

Operations established store 5497 at 1:52 PM with an initial luminosity of 152.89E30.

The MI-62 LCW came into alarm at 3:28 PM on supply high temperature. The Duty Mechanic was called to clean the pond strainers for this system.

Operations reported at 4:24 PM that the Recycler had stashed antiprotons two times during the day shift.

MI-62 pond strainers were cleaned by 5:27 PM.

Operations reported at 11:56 PM that the Recycler had stashed antiprotons four times during the evening shift.

Saturday June 16

The midnight shift began with Operations monitoring store 5497 (54.43E30) and the stash (185.9E10), with Pbar stacking (16.7E10), and with NuMI and MiniBooNE taking beam.

Operations reported at 8:00 AM that the Recycler had stashed antiprotons three times during the midnight shift.

Operations terminated store 5497 at 10:02 AM.

An access was made into F-Sector at 10:29 AM to repair a kicker magnet (F17K2).

D0 is making a short 30 minute access beginning at 11:17 AM.

F-Sector and D0 accesses were completed by 11:39 AM

The TeV went into shot setup at 2:00 PM.

Operations established store 5499 at 4:44 PM with an initial luminosity of 207.70E30.

At 5:15 PM the Pbar stacking rate was cut in half, the problem was traced to a failed kicker magnet. An access will be needed to replace the bad magnet.

The access to replace the bad kicker magnet began at 8:00 PM. No beam to PBar while the magnet is being replaced. NuMI and MiniBooNe will continue to take beam.

Sunday June 17

The midnight shift began with Operations monitoring store 5499 and the stash , with Pbar off for kicker magnet replacement, and with NuMI and MiniBooNE taking beam.

The Debuncher kicker has been replaced, beam returns to PBar at 1:37 AM.

Operations reported at 8:00 AM that the Recycler had stashed antiprotons two times during the midnight shift.

Operations reported at 4:00 PM that the Recycler had stashed antiprotons four times during the day shift.

Operations terminated store 5499 at 8:59 PM.

The TeV went into shot setup at 9:12 PM.

At 10:18 PM a high temperature alarm in the Main Injector caused the secondary pond pump at MI-10 to trip off along with the LCW at the same house. When the LCW pump was turned back on, it caused a bulk corrector supply in the Recycler to trip which resulted in the loss of 144e10 out of the stash.

Monday June 18

The midnight shift began with the TeV in shot setup, and with NuMI and MiniBooNE taking beam.

Operations established store 5500 at 12:31 AM with an initial luminosity of 83.68E30.

The Plans for Monday

The plan for today is to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/18/2007    email Fermilab